The given ordered pairs are solutions of the same Linear Equation. Find N. (3,3) (0,-3) (2,N)
Answer by stanbon(75887)

The given ordered pairs are solutions of the same Linear Equation. Find N. (3,3) (0,-3) (2,N)
---
The y-intercept is -3
The slope is (-3-3)/(0-3) = 2
---
f(x) = 2x-3
---
Therefore f(2) = 4-3 = 1
---
Ans: (2,1)
=============
Cheers,
Stan H.